OpenStudy (anonymous):

Suppose that P(A)=0.41, P(C∣A)=0.005, and P(C′∣A′)=0.008. Find P(A∣C).

OpenStudy (anonymous):

baye's formula right?

OpenStudy (anonymous):

so you know how to use it or should we go through step by step?

OpenStudy (anonymous):

*do you know how to use it?

OpenStudy (anonymous):

Yes, but I'm confused with the C' and A', that part throws me off. I understand it without the primes, but with them it's confusing.

OpenStudy (anonymous):

you want \(P(A|C)\) which is by definition \[\frac{P(A\cap C)}{P(C)}\]

OpenStudy (anonymous):

you need to find \(P(C)\) and \(P(A\cap C)\) the second one is easy, because \[P(A\cap C)=P(A)\times P(C|A)\] and you know both of those numbers

OpenStudy (anonymous):

Okay, so P(A∩C)= .00205

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now you need \(P(C)\) and that takes some work

OpenStudy (anonymous):

great

OpenStudy (anonymous):

not much work though

OpenStudy (anonymous):

Alright, how do i do it

OpenStudy (anonymous):

give me a second i think i can come up with an easy way

OpenStudy (anonymous):

yes, lets make it easy by drawing a venn diagram first off \(P(A')=1-.41=.59\) is a no brainer

OpenStudy (anonymous):

then \[P(A'\cap C')=P(C'|A')\times P(A')=.008\times .59=.00427\]

OpenStudy (anonymous):



OpenStudy (anonymous):

now it should be easy to find \(P(C)\) since it all has to add up to 1 is it clear how i filled out the diagram?

OpenStudy (anonymous):

It does, but does the entire thing have ot add up to one or just the circles?

OpenStudy (anonymous):

the entire thing that is why we needed \(P(A'\cap C')\) it is what is outside the two circles

OpenStudy (anonymous):

i get \(1-.41-.00427 = 0.58573\)

OpenStudy (anonymous):

So C would = .58573, and then you'd take .00205/.58573

OpenStudy (anonymous):

aaannnd it says that answers wrong